OpenStudy (anonymous):

\[\sqrt{3} ( 2 + \sqrt{3}) .. ( 3 + \sqrt{5} )^{2} .. \left(\begin{matrix}2 \\ \sqrt{3}\end{matrix}\right)^{1} .... Does anyone know how \to simplify these ??\]

OpenStudy (anonymous):

I don't think I understand your question...? If you mean\[ \sqrt{3}\left(2+\sqrt{3}\right)=2\sqrt{3}+3 \]And: \[ \left(3+\sqrt{5}\right)^2=\left(3+\sqrt{5}\right)\left(3+\sqrt{5}\right)=3^2+2\cdot3\sqrt{5}+\left(\sqrt{5}\right)^2=\\9+6\sqrt{5}+5=14+6\sqrt{5} \]